OpenStudy (anonymous):

integral of x^2 / (9sqrt3 * sec^5 x)

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ x^2 }{ 9\sqrt{3} \sec^5x }dx\] is this what you have?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I'm going to assume you do. Okay so the integrand (what's inside the integral can rewritten as: \[\frac{ 1 }{ 9\sqrt{3} }\frac{ x^2 }{ \sec^5x }\] and remember when doing integrals when you have a product like above where one is a constant (a number that doesn't involve variables) you can put it in front of the integral so we have: \[\frac{ 1 }{ 9\sqrt{3} }\int\limits_{}^{}\frac{ x^2 }{ \sec^5x }dx\]

OpenStudy (anonymous):

Ok, I get that.

OpenStudy (anonymous):

What's next?

OpenStudy (anonymous):

Next we play around with the sec^5x what can we rewrite it as?

OpenStudy (anonymous):

(sec^2 x)^2 (secx)?

OpenStudy (anonymous):

or, turn it into cos?

OpenStudy (anonymous):

So do you think you can do \[\int\limits_{}^{}x^2\cos^5x dx\]

OpenStudy (anonymous):

since secx=1/cosx

OpenStudy (anonymous):

Then it's x^2 (1-sin^2 x)^2 (cos x)?

OpenStudy (anonymous):

that second factor needs to be to the power of 4

OpenStudy (anonymous):

I have to go but here is a link http://integrals.wolfram.com/index.jsp?expr=%28x%5E2%29*%28cosx%29%5E5&random=false

OpenStudy (goformit100):

Yeah That Too Helped me^

OpenStudy (anonymous):

@goformit100 I'm not sure where to go from there. Can you help me?

OpenStudy (experimentx):

use By parts .. write cos^5(x) as (1-sin^2x)^2 cos(x) to evaluate integral of cos^5(x)

OpenStudy (anonymous):

I got sinx -(2sin^3 x)/3 + sin^5/5 for the integral of cos^5

OpenStudy (experimentx):

use integration by parts again to get rid of x

OpenStudy (anonymous):

I think I'm doing it wrong. I'm about to do integration by parts 3 times.

OpenStudy (experimentx):

you deal with that constant .. i'll write the outline.